# Scaling Reproducibility: An AI-Assisted Workflow

### v1

@@ C. All 215 Specifications Input Success Failure Success rate
Notes. Panels A and B report paper-level success: a paper succeeds at a given stage if at least one specification passes. Panel C reports specification-level success. “Failure” at the material retrieval stage reflects that, among the 67 original papers, replication materials for 12 are no longer publicly available online and were manually supplied from archived copies.
Performance generalizes to the expanded sample. For the 25 newly incorporated papers, all stages succeed, yielding a 100% end-to-end rate. At the specification level, all 215 models are successfully extracted, executed, and analyzed; the 90% overall rate reflects out-of-scope designs rather than execution failures.
At first glance, this result may seem surprising. Three factors account for it. First, the system is version-controlled and adaptive: recurring execution issues are encoded as repair rules and therefore do not recur under a fixed pipeline version. Second, as noted above, the papers in the expanded sample were published after 2023, when all three journals required in-house replication. Third, the diagnostic template is deliberately narrow and well defined. Once the benchmark 2SLS point estimate is reproduced, downstream diagnostics are mechanically determined by the template. Conditional on correct specification parsing and estimation, the remaining steps proceed deterministically.
The workflow substantially reduces time and monetary cost relative to manual replication. For papers with accessible materials, end-to-end processing completes within minutes per paper, ranging from under one minute to half an hour, and can be fully parallelized. Most wall-clock time is spent downloading replication packages and running LLM diagnostics; only a small share of computation requires LLM calls. Once the code is stabilized in the execution layer, the marginal cost per additional paper is low. Researchers remain responsible for interpreting diagnostic reports, but core execution is automated.
5.2 Empirical Results from the Extended IV Corpus
We apply the automated workflow to the extended IV corpus using the same diagnostic template as in Lal et al. (2024). The main empirical patterns, after the AI workflow applies the diagnostic template, remain similar to those reported in the original study.
[…]
These findings reinforce the argument in Lal et al. (2024) that many estimates from observational IV designs rest on fragile identification assumptions, including instrument unconfoundedness and the exclusion restriction. In practice, such designs are often introduced to strengthen the credibility of causal claims based on “naive” OLS; however, when these assumptions fail, especially in the presence of publication bias toward statistically significant results, bias in 2SLS estimates may be substantially larger than in OLS.
A key difference, however, is the time required: whereas the original study involved approximately four years of manual data collection, processing, replication, and reanalysis, the expanded corpus is processed within days under the automated pipeline.
6. Discussion
This paper presents an AI-orchestrated workflow for systematic diagnostic evaluation of empirical research with publicly available data and code. The system executes end to end and generates standardized reports at scale with minimal human intervention. It relies, however, on diagnostic templates defined by human experts. While extraction and execution are automated, benchmark estimands, inclusion criteria, and warning thresholds reflect prior methodological judgment. The workflow therefore scales evaluation rather than defining evaluative standards. Related evidence from Straus and Hall (2026) shows that frontier coding models can reproduce and extend published political science analyses when given structured access to data and code.
Human-led replication initiatives remain indispensable. Journal-based data editor programs (e.g., the AEA Data Editor and Political Analysis replicators), research transparency organizations such as OSF and BITSS, the Institute for Replication (I4R), and independent replication communities continue to play a central role in credibility assessment. Evaluating identification strategies, measurement choices, data construction, and research design requires substantive judgment that cannot, at present, be delegated to automated systems. The workflow proposed here is complementary to these efforts, not a substitute.
We regard this separation between expert judgment and automated execution as a feature. Once templates and standards are specified, the pipeline applies them uniformly across studies, promoting consistency and transparency. As AI systems improve, elements of template design may become more adaptive. Because the architecture is modular, such changes can be incorporated without altering its core structure.
We outline extensions and discuss broader implications for empirical research.
6.1 Future Work
We consider three future extensions: expanding across research designs, deepening the scope of replication, and integrating the workflow into research infrastructure.
Extending to other designs.
A natural extension applies the workflow to additional research designs for which we have previously constructed structured corpora. One candidate is panel studies estimated using estimators tied to two-way fixed effects (TWFE) models under variants of the parallel trends assumption. In earlier work, we assembled and reanalyzed a corpus of 49 such studies (Chiu et al., 2023). There, the benchmark object is the baseline TWFE estimate, and the diagnostic objective is to assess robustness to heterogeneous treatment effects using modern difference-in-differences estimators. Because the corpus was built under explicit inclusion criteria and a clearly defined estimand, it offers a disciplined template for automation. The workflow can replicate this corpus and extend it to newly published studies under the same criteria. A related extension concerns studies that estimate heterogeneous treatment effects using linear interaction models. Hainmueller et al. (2019) compiled a corpus of 22 such applications. Together with the updated diagnostic template in Liu et al. (2025), this corpus provides a structured foundation for scaling.
From diagnostics to full replication.
The current workflow evaluates benchmark specifications. A natural next step is to reconstruct complete tables and figures and verify internal coherence across reported results. This would allow systematic comparison between reported and regenerated outputs. For example, the system could easily detect when two tables claim to use the same sample but report drastically different numbers of observations, or when a robustness specification silently changes clustering levels.
More broadly, the workflow makes previously costly diagnostics more feasible. It can automate cluster jackknife and leave-one-out procedures, compute leverage and Cook’s distance, apply alternative winsorization thresholds, and implement permutation or bootstrap inference under alternative resampling schemes. In most empirical settings, these analyses are technically possible but rarely reported because they require substantial time and computation. By standardizing and parallelizing these procedures, the system lowers the cost of applying them across many studies. Scaling from individual specifications to integrated studies shifts attention from isolated coefficients to the stability of empirical claims as a whole.
Integration with the research process.
The workflow can also integrate into the research and publication pipeline. It can assist authors in preparing replication materials through standardized, machine-readable structures that clarify file organization and specification definitions. Automated validation can identify missing dependencies, unresolved paths, or incomplete documentation at submission rather than after publication.
If journals adopt such tools, reproducibility checks could become more routine. Centralized verification services or third-party replication teams could rely on standardized diagnostic outputs, reducing the burden on editors and shortening review cycles. In this role, the workflow functions not only as a diagnostic instrument but as part of the infrastructure that supports more systematic verification.
6.2 Implications for Empirical Research
We expect this or similar workflows to influence how empirical research is conducted, evaluated, and accumulated in the coming years. We outline several implications below.
Lowering the cost of verification dramatically.
The most immediate implication is a substantial reduction in the marginal cost of verification. When replication materials are available, recomputing estimates and applying standardized diagnostics becomes considerably less expensive than under current practice. This does not resolve disputes about identification or theory, but it changes incentives. At present, systematic verification is rare because its cost often exceeds its expected benefit. As that cost falls, more journals may find it feasible to require in-house or third-party reproducibility checks as a condition for acceptance. Authors, anticipating a higher likelihood of auditing, may adopt more disciplined coding practices and address influence, clustering, and resampling concerns ex ante. Verification thus becomes more closely integrated into the publication process rather than applied only after controversies arise.
Standardizing diagnostic reporting.
Harmonized analysis datasets and structured metadata enable large-scale reanalysis under consistent criteria. Each of our previous large-scale reanalysis projects took three to four years of sustained effort. Much of that time was spent harmonizing replication materials, clarifying benchmark estimands, and standardizing robustness checks across heterogeneous applications. With the present workflow, many of these steps can be automated, substantially reducing the time required to conduct comparable large-scale reanalyses.
Taken together, these implications suggest that agentic AI workflows for reproducibility, with humans in the loop, can function as research infrastructure. They do not replace researchers’ substantive judgment, but make systematic evaluation easier to conduct and harder to avoid. By lowering the cost of verification, standardizing diagnostics, and accelerating methodological development, they may help help make transparency and cumulative scrutiny part of routine empirical practice.
Supplementary Materials
Scaling Reproducibility: An AI-Assisted Workflow
for Large-Scale Reanalysis
  1. A.
    System Architecture and Workflow
    • A.1.
      Three-Layer Architecture
    • A.2.
      Agents and Stage-Level Implementation
  2. B.
    IV Diagnostic Template
  3. C.
    Adaptive Execution Mechanism
    • C.1.
      The Resolution Cycle
    • C.2.
      Adaptation in Practice
  4. D.
    Empirical Inventory and Performance
    • D.1.
      Classes of Implementation Variation
    • D.2.
      Pipeline Performance
    • D.3.
      Detailed Inventory of Resolved Issues
Appendix A System Architecture and Workflow
@@
A.1 Three-Layer Architecture
@@
A.2 Agents and Stage-Level Implementation
This section describes the AI agents that implement each stage of the workflow. The focus is operational: inputs, outputs, implementation details, and design decisions that shaped the system through repeated encounters with real replication packages. Statistical diagnostics are described separately in Section B.
A.2.1 Data Acquisition: Profiler and Librarian
The first step in reproducing an empirical study is obtaining the paper itself and its accompanying replication package—typically comprising data files and analysis code. These tasks are handled by the Profiler and the Librarian. Together they illustrate the adaptive-orchestration-with-deterministic-computation principle at the pipeline’s entry point: the Profiler interprets unstructured PDF text, while the Librarian executes deterministic retrieval routines.
Profiler (Metadata Extraction)
Given a paper in PDF format, Profiler converts the PDF to plain text using pdftotext, then extracts structured metadata including title, authors, publication year, and journal. More critically, it searches for a “Data Availability Statement” (or equivalent section) and extracts the data repository URL—for example, links to Harvard Dataverse, GitHub, or the Open Science Framework (OSF).
Extraction combines deterministic parsing rules (e.g., URL detection, section matching) with limited language-model assistance when formatting is irregular or nonstandard. The extracted information is written to a standardized JSON file (study_info.json), which serves as the input to the Librarian.
Librarian (Replication Package Retrieval)
The Librarian reads the repository URL recorded in study_info.json and downloads the complete replication package. It supports commonly used academic data hosting platforms:
  • Dataverse: retrieves the dataset file list via the Dataverse API and downloads each file individually.
  • GitHub: downloads the repository as a ZIP archive.
  • OSF: retrieves project files via the OSF API.
  • Direct HTTP links: downloads files via standard HTTP requests.
A key design choice is a PDF-first retrieval strategy: URLs embedded in the paper are prioritized over keyword-based search. Title-based search frequently returns irrelevant datasets for papers with common keywords. By contrast, URLs in the PDF typically point directly to the authors’ own repository. Keyword search is used only as a fallback when no URL is found. If automated retrieval fails entirely, the user may manually supply the replication package so that downstream stages can proceed.
The output is a versioned local copy of the full replication package.
A.2.2 Specification Identification: Profiler (Again)
With the replication package available, Profiler (the same agent in charge of metadata extraction) identifies the 2SLS specifications implemented in the code. This is one of the most technically challenging stages of the pipeline.
Multi-Language Code Parsing
The Profiler parses all Stata (.do), R (.R), and Python (.py) scripts in the replication package. It uses language-specific deterministic patterns to detect IV estimation commands.
For Stata, matched commands include ivreg2, ivregress 2sls, reghdfe ... (D = Z), and ivprobit. For R, it detects ivreg(), iv_robust(), and feols() calls containing IV syntax. For Python, it detects IV2SLS() commands.
For each identified IV command, the Profiler extracts a structured representation of the specification, including the outcome variable (YY), endogenous treatment (DD), instrument(s) (ZZ), exogenous controls (XX), clustering variable, fixed effects, statistical software used, source script file, and, when available, the corresponding table reference in the paper. These fields are recorded in metadata.json, which is consumed by the Runner and Skeptic.
Multiple Specification Handling
Empirical papers often report multiple IV specifications. The Profiler extracts all IV commands and then selects primary specifications through a multi-step procedure:
  1. 1.
    Deduplicate specifications based on (Y,D,Z,X)(Y,D,Z,X) combinations.
  2. 2.
    Rank candidates using heuristic rules (e.g., explicit table references, location in main-results scripts, richer control sets).
  3. 3.
    Retain up to three primary specifications per study.
Language Model Assistance
When deterministic parsing fails due to opaque variable names, sparse comments, or nonstandard syntax, the Profiler selectively invokes a lightweight language model to interpret the command structure. For example, given a complex Stata command, the model may identify which variable serves as the endogenous regressor and which as the instrument.
Patterns that initially require LLM interpretation are progressively encoded as deterministic parsing rules. Over time, this reduces reliance on model interpretation and increases reproducibility across standard command formats.
A.2.3 Execution Environment: Janitor and Runner
Janitor (Code Preparation)
Replication packages are typically written under assumptions about directory structure and software environment. Direct execution in a different environment often fails due to path mismatches, unavailable graphics devices, deprecated packages, or syntax conventions.
The Janitor performs automated repairs to enable execution in the current environment. These operations include:
[…]
  • Graphics suppression: Commenting out graphical commands (e.g., graph export, pdf(), png()) that may fail in headless environments.
  • Dependency handling: Substituting or commenting out deprecated R packages (e.g., rgdal, rgeos, maptools).
  • Stata-specific repairs: Handling #delimit syntax, macro substitution (e.g., \$datadir), and panel declarations.
  • Data format conversion: Converting nonstandard formats (e.g., .tab) to formats readable by Stata or R.
Each operation was added in response to a real failure class encountered during development. All modifications are recorded in a structured cleaning log (cleaning_log.json).
Runner (Execution and Data Extraction)
The Runner executes the cleaned replication code in batch mode using the appropriate interpreter: stata -b do, Rscript, or python3, depending on the software type recorded in metadata.json.
The primary objective is to extract the analysis dataset containing all variables used in the reported specification. Data export commands are inserted at appropriate execution points, and the in-memory dataset is saved as CSV. For Stata studies, .dta files are also converted to CSV after execution.
[…]
Outputs include the extracted dataset (analysis_data.csv), execution logs (execution_log.txt), and cross-language validation results.
A.2.4 Diagnostic Execution: Skeptic
The Skeptic applies the human-designed diagnostic template to the extracted dataset by invoking a standalone R script that implements the IV diagnostics described in Section B. The Skeptic contains no language-model calls and no adaptive logic. All statistical procedures are deterministic and version-controlled.
Diagnostic results, including all statistics and warning indicators, are written to a file called diagnostics.json, which serves as the input to the reporting stage.
After computing all diagnostics, the Skeptic assigns a summary credibility rating based on the number of triggered warning flags:
Table S1: Credibility Rating Scheme
Rating Condition Interpretation
High No warnings Strong instruments and robust inference
Moderate 1–2 warnings Some concerns; interpret with caution
Low 3–4 warnings Substantial validity concerns
Very Low 5$\geq$ 5 warnings Results likely unreliable
This rating is an auxiliary summary indicator rather than a substitute for independent evaluation. Individual diagnostics may differ in substantive importance—for example, extremely weak instruments may be more consequential than several minor warnings. The rating is intended to facilitate screening and to highlight specifications that merit closer scrutiny.
A.2.5 Reporting: Journalist
The Journalist converts diagnostics.json into a standardized Markdown report and visualizations. The report includes model specification details, diagnostic statistics, and warning summaries. Standardized charts include first-stage FF-statistics, bootstrap confidence intervals, and jackknife sensitivity distributions.
Because reporting is fully deterministic and template-based, identical inputs produce identical outputs, preserving cross-study comparability.
Appendix B IV Diagnostic Template
@@
Appendix C Adaptive Execution Mechanism
The preceding sections described the self-improvement mechanism in architectural terms. This section makes that mechanism concrete. We document the recurring resolution cycle through which new classes of implementation variation were addressed, and the role of structured knowledge accumulation in preventing recurrence. Representative examples are drawn from the 92-paper evaluation corpus. A complete empirical inventory of resolved issue classes appears in Appendix D.
C.1 The Resolution Cycle
The workflow expands coverage through a recurring four-phase cycle.
First, the system detects anomalies. These include execution failures, silent data inconsistencies, and mismatches between expected and observed outputs. It identifies them by reading execution logs, comparing estimates across software, and checking intermediate files for internal consistency.
Second, it identifies the root cause. Rather than retrying commands, it traces the failure across stages and programming languages to locate the source of the problem.
Third, it implements a general fix. The repair is written as a rule in the relevant deterministic code layer, not as a patch for a single paper. Future papers exhibiting the same pattern are handled automatically.
Fourth, it records the resolution. The context, problem, fix, and impact are documented in the agent’s knowledge base (SKILL.md) in a standardized format so that the solution can be retrieved and reused.
This cycle differs from manual replication and fixed scripts. Manual replication repeats these steps for each paper, and lessons are rarely formalized. A fixed script applies predefined rules but cannot detect new failure classes or extend its rule set. Here, the same four steps are applied systematically, with human review limited to approving code updates before they are committed.
Adaptation proceeds through two channels. When a failure class is well-defined and recurring, the fix is encoded as a deterministic rule in the relevant agent’s code. When a pattern is better captured as contextual guidance, it is recorded in the agent’s knowledge base and used to guide future orchestration. In both cases, updates occur between runs and are version-controlled.
C.2 Adaptation in Practice
We describe every steps of the adaptation cycle below and illustrate them with examples from practice.
C.2.1 Discovery
Many consequential failures produced no explicit error message. Detection therefore relied on cross-field consistency checks.
Example: Silent coefficient failure due to backtick quoting.
When CSV column names begin with an underscore (e.g., _log_providers), R internally wraps them in backticks. A call such as coef(fit)["_log_providers"] returns NA silently. The orchestrator detected this anomaly because first-stage statistics were computed correctly while downstream estimates—2SLS coefficients, jackknife results, bootstrap intervals—were missing. This pattern suggested a name-matching failure rather than a data or model error. The fix introduced a four-way lookup function that attempts raw, backtick-quoted, stripped, and fuzzy matches, and was applied across the diagnostic script.
Example: Destructive safeguard in Stata.
The pattern capture drop X; gen X = expr prevents error r(110) when a variable already exists. However, if expr references a missing variable, gen fails after drop has removed the original variable. The orchestrator identified this when a replication package contained a precomputed variable that was inadvertently deleted. The fix replaced drop-and-recreate with a backup–restore pattern: rename the original variable, attempt generation, and restore if generation fails.
C.2.2 Root-Cause Diagnosis
Complex failures often spanned multiple agents and programming environments.
Example: Cross-language sign reversal.
In one specification, R cross-validation yielded +2.15+2.15 while Stata produced 2.24-2.24. The orchestrator traced the discrepancy across seven layers: a dropped lagged control, variable-name abbreviation in Stata, incomplete panel export, and interaction with e(sample) memory. Switching to a full-panel export with an estimation-sample indicator resolved the discrepancy, producing an exact coefficient match (2.2420-2.2420). Diagnosis required coordinated reasoning across Python, Stata, and R components.
C.2.3 Generalized Code Repair
Fixes were implemented as general rules rather than paper-specific patches.
First, when a single failure revealed a broader pattern, we generalized the rule. For example, failure to recognize abbreviated #delimit commands led to expansion of the regular-expression coverage to include all valid abbreviations. The change applied to all subsequent papers using any shortened form.
Second, when repeated failures exposed a structural weakness, we redesigned the component. Retrieval based on keyword search frequently downloaded incorrect datasets. We therefore adopted a PDF-first retrieval strategy that prioritizes repository links embedded in the paper.
Third, when insertion errors arose during data export, we refined the code-injection logic. Exporting analysis data required detecting the active delimiter mode, scanning multi-line commands before insertion, and using content-based anchors rather than line numbers to avoid drift as earlier edits shifted subsequent lines.
C.2.4 Knowledge Accumulation
Each resolved issue is recorded in structured form:
[Date] capture drop + gen safeguard destroys variables
Context: Pattern capture drop X; gen X = EXPR.
Problem:
gen failure deletes preexisting variable.
Fix: Backup–restore strategy.
Impact: Generalizable to all similar exports.
Entries are standardized, retrievable, and version-controlled. Across agents, the knowledge base expanded to 64 entries, covering syntax handling, name resolution, translation mappings, specification extraction, and workflow coordination.
@@ Appendix D Empirical Inventory and Performance
This section documents the full set of implementation-level issues encountered and resolved during development and evaluation of the AI-assisted reproduction pipeline across 92 IV studies. Each entry records a distinct execution pattern, the agent(s) adjusted, and the corresponding resolution. The detailed inventory is organized into ten classes. Together, they provide a concrete account of the recurring irregularities that arise when replication materials are executed in a standardized, automated environment.
@@ D.1 Classes of Implementation Variation
Table S2: Classes of variation encountered and resolved across 92 papers

### v2

@@ C. All 215 Specifications Input Success Failure Success rate
Notes. Panels A and B report paper-level success: a paper succeeds at a given stage if at least one specification passes. Panel C reports specification-level success. “Failure” at the material retrieval stage reflects that, among the 67 original papers, replication packages for 12 are no longer publicly available online and were manually supplied from archived copies.
Appendix B System Architecture and Workflow
@@
B.1 Three-Layer Architecture
@@
B.2 Agents and Pipeline Stages
The workflow is implemented through a set of specialized agents, each responsible for a distinct stage of the pipeline. The agents serve both the full-paper replication task (Phases A and B) and the design-specific diagnostic task (Phase C). The IV diagnostic template is described separately in Section C.
Phase A: Acquisition and execution.
These tasks are distributed across four agents. The Profiler extracts metadata and repository links from the published PDF, prioritizing URLs embedded in the paper over keyword-based search. The Librarian retrieves the replication package from public repositories (Dataverse, GitHub, OSF, or direct links). The Janitor prepares the code for automated execution: path dependencies are resolved, deprecated packages are handled, and environment assumptions are standardized. The Runner executes the complete codebase and captures all estimation outputs—coefficients, standard errors, sample sizes, and clustering information—in structured form.
For the IV diagnostic task, the workflow additionally identifies the target 2SLS specifications in the code using language-specific parsing across Stata, R, and Python. Up to three primary specifications per study are selected for diagnostic analysis. The analysis dataset for each specification is extracted and, when the original code is in Stata, independently re-estimated in R to verify cross-language consistency.
Phase B: Reproducibility verification.
The Extractor extracts all coefficients reported in the paper’s regression tables using a vision-capable language model that reads rendered table images. The Matcher compares extracted coefficients against code-generated outputs through precision-aware rounding and optimal one-to-one assignment, producing a per-table match rate and an overall reproducibility verdict. Phase B applies only to the full-paper replication task.
Phase C: Diagnostic evaluation.
For papers that pass the reproducibility check and have a supported research design, the Skeptic applies a standardized diagnostic template. For IV designs, this includes first-stage FF-statistics, Anderson–Rubin tests, bootstrap confidence intervals, jackknife sensitivity, and OLS comparisons, implemented through established statistical packages (Lal and Xu, 2024; Blair et al., 2024; Berge, 2023). All diagnostic procedures are deterministic and version-controlled. The Journalist compiles results into a standardized report with specification-level diagnostics and summary indicators.
B.3 Adaptive Execution
Replication packages span multiple programming languages, directory structures, and coding conventions. Many failure modes arise only when new materials are encountered. The workflow addresses this heterogeneity through a structured adaptation mechanism.
When the workflow encounters a recurring failure class—such as path mismatches, deprecated dependencies, or nonstandard syntax—the resolution is encoded as a generalized rule in the deterministic execution layer rather than as a paper-specific patch. Each resolution is documented in a structured knowledge base associated with the relevant pipeline stage, recording the failure context, root cause, fix, and scope of applicability. These knowledge bases are version-controlled: coverage expands across pipeline versions, while numerical behavior remains fixed within each version.
Adaptation proceeds through two channels. Well-defined, recurring failure classes are encoded as deterministic code rules that apply automatically to future papers. Patterns better captured as contextual guidance are recorded in the knowledge base and used by the orchestrator to inform task routing and error recovery. In both cases, updates occur between runs and are subject to human review before being committed. The empirical inventory of resolved issue classes is reported in Section D.
Appendix C IV Diagnostic Template
@@ Appendix D Empirical Inventory and Performance
This section documents the full set of implementation-level issues encountered and resolved during development and evaluation of the AI-assisted reproduction pipeline across 92 IV studies. Each entry records a distinct execution pattern, the agent(s) adjusted, and the corresponding resolution. The detailed inventory is organized into ten classes. Together, they provide a concrete account of the recurring irregularities that arise when replication packages are executed in a standardized, automated environment.
@@ D.1 Classes of Implementation Variation
Table S5: Classes of variation encountered and resolved across 92 IV studies